OpenStudy (anonymous):

Don't know why, but I'm having trouble deriving this equation: 1/sqrt(x)

OpenStudy (lalaly):

derivative?

OpenStudy (anonymous):

Yes, sorry, I meant *differentiating, not deriving.

OpenStudy (lalaly):

ok\[\sqrt x \text{ means } x^{\frac{1}{2}} \]right?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

But it's actually 1/sqrt(x), so x^(-1/2)

OpenStudy (lalaly):

ok so \[\frac{1}{\sqrt x} = \frac{1}{x^{\frac{1}{2}}} = x^{\frac{-1}{2}}\]

OpenStudy (anonymous):

\[\sqrt{x}/(x ^{2}+x)\] I keep ending up with this after attempting to differentiate it.

OpenStudy (lalaly):

\[\large{\frac{d}{dx} x^{-\frac{1}{2}} = -\frac{1}{2}x^{\frac{-3}{2} } }\]

OpenStudy (anonymous):

How did you find that?

OpenStudy (lalaly):

\[\frac{d}{dx} x^m = mx^{m-1}\]

OpenStudy (lalaly):

its called the power rule

OpenStudy (anonymous):

Ah, I see. Well, thanks.

OpenStudy (anonymous):

Makes perfect sense, thanks.